# Recursive variational Gaussian approximation with the Whittle likelihood for linear non-Gaussian state space models

### v1

@@ 3.1 Univariate linear Gaussian SSM
$y_{t}$ $=x_{t}+\epsilon_{t},\quad\epsilon_{t}\sim N(0,\sigma_{\epsilon}^{2}),\quad t=1,\dots,T,$ (18)
$x_{t}$ $=\phi x_{t-1}+\eta_{t},\quad\eta_{t}\sim N(0,\sigma_{\eta}^{2}),\quad t=2,\dots,T,$ (19)
$x_{1}$ $\sim N\left(0,\frac{\sigma_{\eta}^{2}}{1-\phi^{2}}\right).$ (20)
@@
$p(\bm{\theta})=q_{0}(\bm{\theta})=N\left(\begin{bmatrix}0\\ -1\\ -1\end{bmatrix},\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\\ \end{bmatrix}\right).$ (22)
@@ 3.2 Univariate stochastic volatility model
$y_{t}$ $=\sigma_{t}\epsilon_{t},\quad\epsilon_{t}\sim N(0,1),\quad t=1,\dots,T,$ (23)
@@
$x_{t}$ $=\phi x_{t-1}+\eta_{t},\quad\eta_{t}\sim N(0,\sigma_{\eta}^{2}),\quad t=2,\dots,T,$ (25)
$x_{1}$ $\sim N\left(0,\frac{\sigma_{\eta}^{2}}{1-\phi^{2}}\right).$ (26)
@@
This model can be “linearised” with a log-squared transformation (Ruiz,, 1994),
@@
$p(\bm{\theta})=q_{0}(\bm{\theta})=N\left(\begin{bmatrix}2\\ -3\end{bmatrix},\begin{bmatrix}0.5&0\\ 0&0.5\\ \end{bmatrix}\right),$ (31)
which leads to 95% probability intervals of (0.77, 0.99) for $\phi$, and (0.14, 0.36) for $\sigma_{\eta}$.
@@ 3.3 Bivariate SV model
$\mathbf{y}_{t}$ $=\mathbf{V}_{t}\bm{\varepsilon}_{t},\quad\bm{\varepsilon}_{t}\sim N(\mathbf{0},\mathbf{I}_{2}),\quad t=1,\dots,T,$ (32)
$\mathbf{x}_{t}$ $=\mathbf{\Phi}\mathbf{x}_{t-1}+\bm{\eta}_{t},\quad\bm{\eta}_{t}\sim N(\mathbf{0},\mathbf{\Sigma}_{\eta}),\quad t=2,\dots,T,$ (33)
$\mathbf{x}_{1}$ $\sim N(\mathbf{0},\mathbf{\Sigma}_{\eta_{1}}),$ (34)
@@
$p(\bm{\theta})=q_{0}(\bm{\theta})=N\left(\begin{bmatrix}2\\ 2\\ -2\\ -3\\ 0\end{bmatrix},\begin{bmatrix}0.5&0&0&0&0\\ 0&0.5&0&0&0\\ 0&0&0.5&0&0\\ 0&0&0&0.05&0\\ 0&0&0&0&0.05\\ \end{bmatrix}\right).$ (38)
This leads to 95% probability intervals of $(0.52,0.99)$ for $\Phi_{11}$ and $\Phi_{22}$, $(0.001,0.286)$ for $\Sigma_{\eta_{11}}$, $(-0.101,0.101)$ for $\Sigma_{\eta_{21}}$, and $(0.002,0.261)$ for $\Sigma_{\eta_{22}}$.
@@
3.4 Real data examples
@@ Acknowledgements
S1 The R-VGA and R-VGAL algorithms
In this section, we provide a sketch of the derivations for the R-VGA algorithm of Lambert et al., (2022), which serves as the basis for both the R-VGAL algorithm of Vu et al., (2024) and our R-VGA-Whittle algorithm. This section closely follows Section 2.2.2 in Vu et al., (2024).

### v2

@@ 3.1 Univariate linear Gaussian SSM
$y_{t}$ $=x_{t}+\epsilon_{t},\quad\epsilon_{t}\sim\mathrm{Gau}(0,\sigma_{\epsilon}^{2}),\quad t=1,\dots,T,$ (18)
$x_{t}$ $=\phi x_{t-1}+\eta_{t},\quad\eta_{t}\sim\mathrm{Gau}(0,\sigma_{\eta}^{2}),\quad t=2,\dots,T,$ (19)
$x_{1}$ $\sim\mathrm{Gau}\left(0,\frac{\sigma_{\eta}^{2}}{1-\phi^{2}}\right).$ (20)
@@
$p(\bm{\theta})=q_{0}(\bm{\theta})=\mathrm{Gau}\left(\begin{bmatrix}0\\ -1\\ -1\end{bmatrix},\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\\ \end{bmatrix}\right).$ (22)
@@ 3.2 Univariate stochastic volatility model
$y_{t}$ $=\sigma_{t}\epsilon_{t},\quad\epsilon_{t}\sim\mathrm{Gau}(0,1),\quad t=1,\dots,T,$ (23)
@@
$x_{t}$ $=\phi x_{t-1}+\eta_{t},\quad\eta_{t}\sim\mathrm{Gau}(0,\sigma_{\eta}^{2}),\quad t=2,\dots,T,$ (25)
$x_{1}$ $\sim\mathrm{Gau}\left(0,\frac{\sigma_{\eta}^{2}}{1-\phi^{2}}\right).$ (26)
@@
Following Ruiz, (1994), we apply a log-squared transformation,
@@
$p(\bm{\theta})=q_{0}(\bm{\theta})=\mathrm{Gau}\left(\begin{bmatrix}2\\ -3\end{bmatrix},\begin{bmatrix}0.5&0\\ 0&0.5\\ \end{bmatrix}\right),$ (31)
which leads to 95% probability intervals of $(0.57,0.99)$ for $\phi$, and $(0.11,0.44)$ for $\sigma_{\eta}$. The interval for $\phi$ aligns closely with those used in similar studies of the stochastic volatility (SV) model. For example, Kim et al., (1998) employed the transformation $\phi=2\phi^{*}-1$ with $\phi^{*}\sim\textrm{Beta}(20,1.5)$, resulting in a 95% probability interval for $\phi$ of $(0.59,0.99)$. Additionally, Kim et al., (1998) utilized an inverse gamma prior for $\sigma_{\eta}^{2}$ with shape and rate parameters of $2.5$ and $0.025$, respectively, corresponding to a 95% probability interval $\sigma_{\eta}$ for $(0.06,0.24)$, which is comparable to the interval implied by our chosen prior.
@@ 3.3 Bivariate SV model
$\mathbf{y}_{t}$ $=\mathbf{V}_{t}\bm{\varepsilon}_{t},\quad\bm{\varepsilon}_{t}\sim\mathrm{Gau}(\mathbf{0},\mathbf{I}_{2}),\quad t=1,\dots,T,$ (32)
$\mathbf{x}_{t}$ $=\mathbf{\Phi}\mathbf{x}_{t-1}+\bm{\eta}_{t},\quad\bm{\eta}_{t}\sim\mathrm{Gau}(\mathbf{0},\mathbf{\Sigma}_{\eta}),\quad t=2,\dots,T,$ (33)
$\mathbf{x}_{1}$ $\sim\mathrm{Gau}(\mathbf{0},\mathbf{\Sigma}_{\eta_{1}}),$ (34)
@@
$p(\bm{\theta})=q_{0}(\bm{\theta})=\mathrm{Gau}\left(\begin{bmatrix}2\\ 2\\ -2\\ -3\\ 0\end{bmatrix},\begin{bmatrix}0.5&0&0&0&0\\ 0&0.5&0&0&0\\ 0&0&0.5&0&0\\ 0&0&0&0.05&0\\ 0&0&0&0&0.05\\ \end{bmatrix}\right).$ (38)
For $\Phi_{11}$ and $\Phi_{22}$, this prior leads to 95% probability intervals of $(0.52,0.99)$. For $\Sigma_{\eta_{11}}$, $\Sigma_{\eta_{21}}$, and $\Sigma_{\eta_{22}}$, the 95% probability intervals are $(0.001,0.286)$, $(-0.101,0.101)$, and $(0.002,0.261)$, respectively.
@@
3.4 Real univariate and bivariate examples
@@ Acknowledgements
S1 The R-VGA algorithm
In this section, we provide a sketch of the derivations for the R-VGA algorithm of Lambert et al., (2022), which serves as the basis for our R-VGA-Whittle algorithm.